grain silo has the shape of a right circular cylinder with a half sphere on top. It is used to store various grains (corn, wheat, etc.) sold and delivered to market. The surface area, in square feet, of a grain silo of height h and radius r, both in feet and including the area of the base, is given by the polynomial . How much surface area does a silo have if it has a radius of 30 ft and a height of 40 ft? Use 3.14 for and round to the nearest tenth of a square foot.
Help please?

Answer by nerdybill(7384)

grain silo has the shape of a right circular cylinder with a half sphere on top. It is used to store various grains (corn, wheat, etc.) sold and delivered to market. The surface area, in square feet, of a grain silo of height h and radius r, both in feet and including the area of the base, is given by the polynomial . How much surface area does a silo have if it has a radius of 30 ft and a height of 40 ft? Use 3.14 for and round to the nearest tenth of a square foot.
.
Surface of base:(pi)r^2
Surface of cylinder:2(pi)rh
Surface of half sphere: (1/2)(4)(pi)r^2
.
where
r is radius
h is height
pi is 3.14
.
Surface area = (pi)r^2 + 2(pi)rh + 2(pi)r^2
Surface area = (pi)r(r + 2h + 2r)
.
Plugging in our values:
Surface area = (pi)r(r + 2h + 2r)
Surface area = (3.14)(30)(30 + 2(40) + 2(30))
Surface area = (3.14)(30)(30 + 80 + 60)
Surface area = (3.14)(30)(170)
Surface area = 16014.0 square feet